i need help on this please ASAP
Mrs. Garcia is 1 year less than 3 times as old as he daughter. Seven years from now she will be 4 more tahn twice as old as her daughter will be then. Find both of their ages
Thank you

Answer by stanbon(75887)

Mrs. Garcia is 1 year less than 3 times as old as her daughter. Seven years from now she will be 4 more tahn twice as old as her daughter will be then. Find both of their ages
--------------
Equations:
g = 3d-1
(g+7) = 2(d+7)+4
-------------------------
Substitute for "g" and solve for "d":
3d-1+7 = 2d+14+4
3d+6 = 2d+18
d = 12 yrs (daughter's age now)
----
Solve for "g":
g = 3d-1
g = 3*12-1
g = 35 yrs (Mrs. Garcia's age now)
=======================================
Cheers,
Stan H.
=============